OpenStudy (anonymous):

The length of a diagonal of a square sandbox is 9 feet. What is the approximate length of one side of the sandbox? A:4.5 B:5.2 C:6.4 D:12.7

OpenStudy (anonymous):

I keep getting 1.41 which is weird because I got that exact answer on 2 other questions.

OpenStudy (wolf1728):

If the diagonal = 9 then by the Pythagorean Theorem side² + side ² = 9² 2 * side² = 81 side² = 81 / 2 side² = 40.5 side = sqrt(40.5) side = 6.3639610307

OpenStudy (anonymous):

Thank you! I see what I was doing wrong now!!